# SparX: A Sparse Cross-Layer Connection Mechanism for Hierarchical Vision Mamba and Transformer Networks

### v1

@@ Abstract
Due to the capability of dynamic state space models (SSMs) in capturing long-range dependencies with near-linear computational complexity, Mamba has shown notable performance in NLP tasks. This has inspired the rapid development of Mamba-based vision models, resulting in promising results in visual recognition tasks. However, such models are not capable of distilling features across layers through feature aggregation, interaction, and selection. Moreover, existing cross-layer feature aggregation methods designed for CNNs or ViTs are not practical in Mamba-based models due to high computational costs. Therefore, this paper aims to introduce an efficient cross-layer feature aggregation mechanism for Mamba-based vision backbone networks. Inspired by the Retinal Ganglion Cells (RGCs) in the human visual system, we propose a new sparse cross-layer connection mechanism termed SparX to effectively improve cross-layer feature interaction and reuse. Specifically, we build two different types of network layers: ganglion layers and normal layers. The former has higher connectivity and complexity, enabling multi-layer feature aggregation and interaction in an input-dependent manner. In contrast, the latter has lower connectivity and complexity. By interleaving these two types of layers, we design a new vision backbone network with sparsely cross-connected layers, achieving an excellent trade-off among model size, computational cost, memory cost, and accuracy in comparison to its counterparts. For instance, with fewer parameters, SparX-Mamba-T improves the top-1 accuracy of VMamba-T from 82.5% to 83.5%, while SparX-Swin-T achieves a 1.3% increase in top-1 accuracy compared to Swin-T. Extensive experimental results demonstrate that our new connection mechanism possesses both superior performance and generalization capabilities on various vision tasks. Code will be publicly available at https://github.com/LMMMEng/SparX.
@@ 1 Introduction
Empowered by dynamic State Space Models (SSMs) capable of modeling long-range dependencies with near-linear complexity, Mamba (Gu and Dao 2024) has demonstrated promising performance in natural language processing (NLP) tasks. Building upon this success, researchers have proposed Mamba-based vision models (Zhu et al. 2024; Yang et al. 2024; Liu et al. 2024; Huang et al. 2024; Behrouz, Santacatterina, and Zabih 2024; Pei, Huang, and Xu 2024) to achieve notable results in visual tasks. Such work has made meaningful attempts to introduce linear-time long-range modeling capability into vision models by leveraging SSMs, which are primarily used to capture spatial contextual information from tokens within the same layer. Nonetheless, there might exist both complementarity and redundancy among token features in different layers as they capture image characteristics and semantics at different levels and granularities. Therefore, features across different layers need to be distilled to extract useful information and remove redundancy through aggregation, interaction, and selection. However, existing Mamba-based vision models do not possess such an ability to conduct feature distillation across layers, hindering the realization of the full power of SSMs.
@@
By hierarchically interleaving ganglion and normal layers in a feedforward network, we design a novel Mamba-based vision backbone, namely, Vision Mamba with sparsely cross-connected layers (SparX-Mamba). In addition to inherent advantages of Mamba-based models, our SparX-Mamba achieves an excellent trade-off between performance and computational cost. Notably, our SparX-Mamba differs from conventional vision models that are constructed by simply stacking the same layer design, resulting in similar feature extraction capabilities at every layer. In contrast, the two types of functionally distinct layers of SparX-Mamba facilitate a model to extract more diverse feature representations, thereby enhancing performance. It is worth noting that our SparX mechanism has also been successfully applied to hierarchical vision Transformers, such as Swin (Liu et al. 2021), and the resulting models also have the aforementioned advantages over the original Transformer models.
As shown in Figure 1, the proposed SparX-Mamba demonstrates superior performance over recently proposed VMamba as well as classical CNN and Transformer-based models, with a smaller number of parameters (Params). For instance, when compared with VMamba-T, SparX-Mamba-T achieves a top-1 accuracy of 83.5% on ImageNet-1K, remarkably surpassing the accuracy of VMamba-T (82.5%). When integrated with UperNet for semantic segmentation on the ADE20K dataset, SparX-Mamba-T outperforms VMamba-T by 1.7% in mIoU. Furthermore, when integrated into Mask R-CNN for object detection on the COCO 2017 dataset, SparX-Mamba-T exceeds VMamba-T by 1.3% in APb and even outperforms VMamba-S, whose complexity is nearly 1.5 times that of SparX-Mamba-T. Meanwhile, both the small and base versions of SparX-Mamba also exhibit notable performance improvements over VMamba. Furthermore, as shown in Table 8, our Sparx-Swin-T achieves a 1.3% improvement in top-1 accuracy compared to Swin-T.
@@ 2 Related Work
Vision Transformers. To adapt Transformers from NLP tasks to computer vision tasks, a vision Transformer (ViT) (Dosovitskiy et al. 2021) splits an image into visual tokens through patch embedding, thereby enabling multi-head self-attention (MHSA) to learn token-to-token dependencies. To further generate hierarchical feature representations and improve model efficiency, many subsequent works have adopted pyramidal architecture designs, including window attention (Liu et al. 2021; Dong et al. 2022; Pan et al. 2023), sparse attention (Yang et al. 2021; Wang et al. 2022; Ren et al. 2022; Wu et al. 2023), and conv-attention hybrid models (Dai et al. 2021; Xiao et al. 2021; Li et al. 2022, 2023; Yu et al. 2024a).
Vision Mamba. Since Mamba (Gu and Dao 2024) has achieved outstanding performance in NLP tasks, many researchers have transferred Mamba to computer vision tasks. As the core of Mamba, SSM can model long-range dependencies with near-linear complexity and has shown excellent performance in vision tasks. For instance, ViM (Zhu et al. 2024) introduces a bidirectional SSM module and constructs an isotropic architecture like ViT (Dosovitskiy et al. 2021). Likewise, PlainMamba (Yang et al. 2024) also builds an isotropic architecture with continuous 2D scanning. VMamba (Liu et al. 2024) extends the scanning order to include four directions and is an early SSM-based hierarchical architecture. Subsequently, a series of hierarchical vision Mamba models have been proposed, including MambaMixer (Behrouz, Santacatterina, and Zabih 2024), LocalMamba (Huang et al. 2024), and EfficientVMamba (Pei, Huang, and Xu 2024).
@@ 3.1 Sparse Cross-Layer Connections
Overview. Inspired by how Retinal Ganglion Cells (RGCs) function in the human visual system, we introduce a novel sparse cross-layer connection mechanism named SparX. The goal is efficiently modeling cross-layer communications, generating diverse information flows, and improving feature reuse within hierarchical vision architectures. In this section, we use hierarchical vision Mamba as our reference network architecture, however, our method is also applicable to hierarchical vision Transformers. As shown in Figure 2, SparX has three building components: Dynamic Position Encoding (DPE) (Chu et al. 2022), Mamba block, and the newly-introduced Dynamic Multi-layer Channel Aggregator (DMCA). Among them, DPE leverages a residual 3$\times$3 depthwise convolution (DWConv), which finds widespread adoption in modern vision backbone networks (Li et al. 2023; Chu et al. 2021; Guo et al. 2022). The Mamba block effectively captures long-range dependencies with a near-linear computational complexity. However, the vanilla Mamba block (Gu and Dao 2024) is unsuitable for being directly embedded into vision backbones. In this regard, we employ the VSS block, which has demonstrated promising performance since initially introduced in VMamba (Liu et al. 2024). The proposed DMCA dynamically aggregates, selects, and encodes cross-layer features, giving rise to powerful and robust feature representations. On top of these basic components, two types of network layers are constructed in SparX: ganglion layers (‘DPE $\rightarrow$ DMCA $\rightarrow$ VSS Block’) and normal layers (‘DPE $\rightarrow$ VSS Block’). Normal layers and ganglion layers are interleaved in a feedforward network to enable effective modeling of cross-layer feature interactions in a sparse manner. The basic components and our overall network architecture will be elaborated below.
@@
Sparse Ganglion Layers aim to designate a subset of evenly spaced layers as ganglion layers while all remaining layers are normal layers. To control the density of ganglion layers, we define a hyperparameter called stride (S), which is one plus the number of normal layers between two nearest ganglion layers. We further define two types of cross-layer connections: (1) Intra-connection between a ganglion layer and a normal layer, (2) Inter-connection between two ganglion layers. To set up sparse cross-layer connections, a ganglion layer only has intra-connections with the normal layers between the closest preceding ganglion layer and itself while it builds inter-connections with multiple preceding ganglion layers. The rationale behind such cross-layer connections is that a ganglion layer can be regarded as an information exchange hub, which gathers information from the closest normal layers and exchanges it with other ganglion layers. To ensure that the final output of a network or stage contains rich semantic information, the last layer of a network or stage is typically a ganglion layer. As an example, if we have an 8-layer network and set the stride (S) to 2, the indices of the normal layers are $\left\{1,3,5,7\right\}$ while the indices of the ganglion layers are $\left\{2,4,6,8\right\}$.
@@
Cross-layer Sliding Window is proposed to further improve computational efficiency, inspired by spatial sliding windows. The motivation behind this design is that, despite using the aforementioned sparse connections, deeper networks may still incur high memory costs due to the need to store and access a number of earlier feature maps. To this end, we introduce another hyperparameter M, which restricts a ganglion layer to have inter-connections with the M closest preceding ganglion layers only. According to these two new rules, even without direct connections, information can still flow from shallower layers to deeper layers quickly through a relatively small number of intra-connections and inter-connections. Figure 2 illustrates an example of SparX with S=2 and M=2.
Dynamic Multi-layer Channel Aggregator (DMCA). To selectively retrieve complementary features from previous layers and dynamically model multi-layer interactions, we propose an efficient Dynamic Multi-layer Channel Aggregator (DMCA). As depicted in Figure 3, let $\mathbf{X}\in\mathbb{R}^{C\times N}$ represents the features of a ganglion layer, and $\left\{\mathbf{Y_{1},Y_{2},...,Y_{L}}\right\}\in\mathbb{R}^{C\times N}$ denote features from preceding ganglion and normal layers, where C and N represent the channel and spatial dimensions, respectively. First, we concatenate the features from preceding layers and project the channel dimension to 2C using a linear layer. Then, we evenly split the projected features into two segments along the channel dimension, $\mathbf{Y_{k}}$ and $\mathbf{Y_{v}}$. Afterwards, we design a grouped channel-wise cross-attention (GCCA) mechanism, where query, key, and value refer to $\mathbf{X}$, $\mathbf{Y_{k}}$ and $\mathbf{Y_{v}}$, respectively. Specifically, we divide the channels into multiple groups, and channel-wise cross-attention is computed within each group. The number of groups (G) is always set to 4 in our experiments. Consequently, in our GCCA, the size of the overall attention map $\mathbf{A}$ for all groups is $G\times\frac{C}{G}\times\frac{C}{G}$, which avoids high computational and memory overhead when large spatial resolutions are used.
@@
$\centering\begin{gathered}\mathbf{Y_{k}},\mathbf{Y_{v}}=\mathrm{Split}(W_{1}(\mathrm{Cat}(\mathbf{Y^{\prime}_{1}},\mathbf{Y^{\prime}_{1}},...,\mathbf{Y^{\prime}_{L}})))\\ \mathrm{Q}=W_{2}(R_{1}(\mathbf{X})),\mathrm{K}=W_{3}(R_{2}(\mathbf{Y_{k}})),\mathrm{V}=W_{4}(\mathbf{Y_{v}})\\ \mathbf{Z}=\frac{\mathrm{{Softmax}}(\mathrm{Q}\mathrm{K}^{\mathrm{T}})}{\sqrt{N/r}}\mathrm{V}\end{gathered}\@add@centering$ (1)
@@
Overall, SparX introduces two flexible hyperparameters, namely S and M, making it easier and more efficient to extend the network deeper and wider. In addition, the incorporation of DMCA enables strong cross-layer feature interactions, giving rise to more powerful and robust representations.
@@
3.2 Network Architecture
Based on the proposed sparse cross-layer connections, we have developed a new Mamba-based vision backbone named SparX-Mamba, a 4-stage hierarchical architecture with three different variants: tiny, small, and base. In stage 1, we do not include any ganglion layers due to the high computational overhead when large-resolution inputs are used. That is, stage 1 contains only two normal layers. In stage 2, we designate the second layer as a ganglion layer, which has an intra-connection with its preceding layer. Since Stage 3 has more layers, we set different hyperparameters for the three variants. As listed in Table 1, due to the relatively small depth of the tiny model, we set S to 2 while S = 3 in both small and base models to avoid high computational cost in deeper models. In stage 4, as an exception, we set all layers as ganglion layers since this stage processes the lowest resolution features, resulting in a relatively small computational cost. Meanwhile, to facilitate communications across different stages, the first ganglion layer of each stage connects to a downsampled version of the final feature from the preceding stage. Furthermore, a cross-layer sliding window is only applied within each stage, to prevent high memory consumption due to excessive feature storage.
@@ 3.2 Network Architecture
SparX-Mamba-T 5.2 27 83.5

### v2

@@ Abstract
Due to the capability of dynamic state space models (SSMs) in capturing long-range dependencies with linear-time computational complexity, Mamba has shown notable performance in NLP tasks. This has inspired the rapid development of Mamba-based vision models, resulting in promising results in visual recognition tasks. However, such models are not capable of distilling features across layers through feature aggregation, interaction, and selection. Moreover, existing cross-layer feature aggregation methods designed for CNNs or ViTs are not practical in Mamba-based models due to high computational costs. Therefore, this paper aims to introduce an efficient cross-layer feature aggregation mechanism for vision backbone networks. Inspired by the Retinal Ganglion Cells (RGCs) in the human visual system, we propose a new sparse cross-layer connection mechanism termed SparX to effectively improve cross-layer feature interaction and reuse. Specifically, we build two different types of network layers: ganglion layers and normal layers. The former has higher connectivity and complexity, enabling multi-layer feature aggregation and interaction in an input-dependent manner. In contrast, the latter has lower connectivity and complexity. By interleaving these two types of layers, we design a new family of vision backbone networks with sparsely cross-connected layers, achieving an excellent trade-off among model size, computational cost, memory cost, and accuracy in comparison to its counterparts. For instance, with fewer parameters, SparX-Mamba-T improves the top-1 accuracy of VMamba-T from 82.5% to 83.5%, while SparX-Swin-T achieves a 1.3% increase in top-1 accuracy compared to Swin-T. Extensive experimental results demonstrate that our new connection mechanism possesses both superior performance and generalization capabilities on various vision tasks. Code is publicly available at https://github.com/LMMMEng/SparX.
@@ 1 Introduction
Empowered by dynamic State Space Models (SSMs) capable of modeling long-range dependencies with near-linear complexity, Mamba (Gu and Dao 2024) has demonstrated promising performance in natural language processing (NLP) tasks. Building upon this success, researchers have proposed Mamba-based models for vision tasks (Zhu et al. 2024; Yang et al. 2024; Liu et al. 2024; Huang et al. 2024; Behrouz, Santacatterina, and Zabih 2024; Pei, Huang, and Xu 2024). Such work has made successful attempts to introduce linear-time long-range modeling capability into vision models by leveraging SSMs, which are primarily used to capture spatial contextual information within the same layer. Nonetheless, there might exist both complementarity and redundancy among token features in different layers as they capture image characteristics and semantics at different levels and granularities. Thus features across different layers need to be distilled to extract useful information and remove redundancy through aggregation, interaction, and selection. However, existing Mamba-based vision models do not possess such an ability to conduct feature distillation across layers, hindering the realization of the full power of SSMs.
@@
By hierarchically interleaving ganglion and normal layers in a feedforward network, we design a novel Mamba-based vision backbone, namely, Vision Mamba with sparsely cross-connected layers (SparX-Mamba). In addition to inherent advantages of Mamba-based models, our SparX-Mamba achieves an excellent trade-off between performance and computational cost. Notably, our SparX-Mamba differs from conventional vision models that are constructed by simply stacking the same layer design, resulting in similar feature extraction capabilities at every layer. In contrast, the two types of functionally distinct layers of SparX-Mamba facilitate a model to extract more diverse feature representations, thereby enhancing performance. It is worth noting that our SparX has also been successfully applied to hierarchical vision Transformers, such as Swin (Liu et al. 2021), and the resulting models also have the aforementioned advantages over the original Transformer models. For example, our SparX-Swin-T achieves a 1.3% improvement in top-1 accuracy compared to Swin-T.
As shown in Figure 1, the proposed SparX-Mamba demonstrates superior performance over recently proposed VMamba as well as classical CNN and Transformer-based models, with a smaller number of parameters (Params). For instance, when compared with VMamba-T, SparX-Mamba-T achieves a top-1 accuracy of 83.5% on ImageNet-1K, remarkably surpassing the accuracy of VMamba-T (82.5%). When integrated with UperNet for semantic segmentation on the ADE20K dataset, SparX-Mamba-T outperforms VMamba-T by 1.7% in mIoU. Furthermore, when integrated into Mask R-CNN for object detection on the COCO 2017 dataset, SparX-Mamba-T exceeds VMamba-T by 1.3% in APb and even outperforms VMamba-S, whose complexity is nearly 1.5 times that of SparX-Mamba-T. Meanwhile, both the small and base versions of SparX-Mamba also exhibit notable performance improvements over VMamba.
@@ 2 Related Work
Vision Transformers. To adapt Transformers from NLP tasks to computer vision tasks, a vision Transformer (ViT) (Dosovitskiy et al. 2021) splits an image into visual tokens through patch embedding, thereby enabling multi-head self-attention (MHSA) to learn token-to-token dependencies. To further generate hierarchical feature representations and improve model efficiency, many subsequent works have adopted pyramidal architecture designs, including window attention (Liu et al. 2021; Dong et al. 2022; Pan et al. 2023), sparse attention (Yang et al. 2021; Wang et al. 2022; Ren et al. 2022; Wu et al. 2023), and conv-attention hybrid models (Dai et al. 2021; Xiao et al. 2021; Li et al. 2022, 2023; Lou et al. 2023; Yu et al. 2024a).
Vision Mamba. Since Mamba (Gu and Dao 2024) has achieved outstanding performance in NLP tasks, many researchers have transferred Mamba to computer vision tasks. As the core of Mamba, SSM can model long-range dependencies with near-linear complexity and has shown excellent performance in vision tasks. For instance, ViM (Zhu et al. 2024) introduces a bidirectional SSM module and constructs an isotropic architecture like ViT (Dosovitskiy et al. 2021). Likewise, PlainMamba (Yang et al. 2024) also builds an isotropic architecture with continuous 2D scanning. VMamba (Liu et al. 2024) extends the scanning order to include four directions and is an early SSM-based hierarchical architecture. Subsequently, a series of hierarchical vision Mamba models have been proposed, including MambaMixer (Behrouz, Santacatterina, and Zabih 2024), LocalMamba (Huang et al. 2024), EfficientVMamba (Pei, Huang, and Xu 2024), and MSVMamba (Shi, Dong, and Xu 2024).
@@ 3.1 Sparse Cross-Layer Connections
Overview. Inspired by how Retinal Ganglion Cells (RGCs) function in the human visual system, we introduce a novel sparse cross-layer connection mechanism named SparX. The goal is efficiently modeling cross-layer communications, generating diverse information flows, and improving feature reuse within Mamba-based architectures. As shown in Figure 2, SparX has three building components: Dynamic Position Encoding (DPE) (Chu et al. 2022), Mamba block, and the newly-introduced Dynamic Multi-layer Channel Aggregator (DMCA). Among them, DPE leverages a residual 3$\times$3 depthwise convolution (DWConv), which finds widespread adoption in modern vision backbone networks (Li et al. 2023; Chu et al. 2021; Guo et al. 2022). The Mamba block effectively captures long-range dependencies with a near-linear computational complexity. However, the vanilla Mamba block (Gu and Dao 2024) is unsuitable for being directly embedded into vision backbones. In this regard, we utilize the VSS block, which has shown promising performance since its introduction in VMamba (Liu et al. 2024). To enhance channel mixing, a ConvFFN (Wang et al. 2022) is incorporated within VSS following (Shi, Dong, and Xu 2024). The proposed DMCA dynamically aggregates, selects, and encodes cross-layer features, giving rise to powerful and robust feature representations. On top of these basic components, two types of network layers are constructed in SparX: ganglion layers (“DPE $\rightarrow$ DMCA $\rightarrow$ VSS Block”) and normal layers (“DPE $\rightarrow$ VSS Block”). These two layers are interleaved in a feedforward network to enable effective modeling of cross-layer feature interactions in a sparse manner.
@@
Sparse Ganglion Layers aim to designate a subset of evenly spaced layers as ganglion layers while all remaining layers are normal layers. To control the density of ganglion layers, we define a hyperparameter called stride (S), which is one plus the number of normal layers between two nearest ganglion layers. We further define two types of cross-layer connections: (1) Intra-connection between a ganglion layer and a normal layer; (2) Inter-connection between two ganglion layers. To set up sparse cross-layer connections, a ganglion layer only has intra-connections with the normal layers between the closest preceding ganglion layer and itself while it builds inter-connections with multiple preceding ganglion layers. The rationale behind such cross-layer connections is that a ganglion layer can be regarded as an information exchange hub, which gathers information from the closest normal layers and exchanges it with other ganglion layers. To ensure that the final output of a network or stage contains rich semantic information, the last layer of a network or stage is typically a ganglion layer. As an example, if we have an 8-layer network and set the stride (S) to 2, the indices of the normal layers are $\left\{1,3,5,7\right\}$ while the indices of the ganglion layers are $\left\{2,4,6,8\right\}$.
Cross-layer Sliding Window is proposed to further improve computational efficiency, inspired by spatial sliding windows. The motivation behind this design is that, despite using the aforementioned sparse connections, deeper networks may still incur high memory costs due to the need to store and access a number of earlier feature maps. To this end, we introduce another hyperparameter M, which restricts a ganglion layer to have inter-connections with the M closest preceding ganglion layers only. Based on these two new rules, even without direct connections, information can still flow from shallower layers to deeper layers quickly through a relatively small number of intra-connections and inter-connections. Figure 2 illustrates an example of SparX with S=2 and M=2.
Dynamic Multi-layer Channel Aggregator (DMCA). To selectively retrieve complementary features from previous layers and dynamically model multi-layer interactions, we propose an efficient Dynamic Multi-layer Channel Aggregator (DMCA). As depicted in Figure 3, let $\mathbf{X}\in\mathbb{R}^{C\times N}$ represents the features of a ganglion layer, and $\left\{\mathbf{Y_{1},Y_{2},\cdots,Y_{L}}\right\}\in\mathbb{R}^{C\times N}$ denote features from preceding ganglion and normal layers, where C and N represent the channel and spatial dimensions, respectively, while L denotes the total number of features from prior layers. First, we concatenate the features from preceding layers and project the channel dimension to 2C using a linear layer. Then, we evenly split the projected features into two segments along the channel dimension, $\mathbf{Y_{k}}$ and $\mathbf{Y_{v}}$. Afterwards, we design a cross-layer grouped channel attention (CGCA) mechanism, where query, key, and value refer to $\mathbf{X}$, $\mathbf{Y_{k}}$, and $\mathbf{Y_{v}}$, respectively. Specifically, we divide the channels into multiple groups, and channel-wise cross-attention is computed within each group. The number of groups (G) is always set to 4 in our experiments. Consequently, in our CGCA, the size of the overall attention map $\mathbf{A}$ for all groups is $G\times\frac{C}{G}\times\frac{C}{G}$, which avoids high computational and memory overhead when large spatial resolutions are used.
@@
$\centering\begin{gathered}\mathbf{Y_{k}},\mathbf{Y_{v}}=\mathrm{Split}(W_{1}(\mathrm{Cat}(\mathbf{Y_{1}},\mathbf{Y_{1}},\cdots,\mathbf{Y_{L}})))\\ \mathrm{Q}=W_{2}(R_{1}(\mathbf{X})),\mathrm{K}=W_{3}(R_{2}(\mathbf{Y_{k}})),\mathrm{V}=W_{4}(\mathbf{Y_{v}})\\ \mathbf{Z}=\frac{\mathrm{{Softmax}}(\mathrm{Q}\mathrm{K}^{\mathrm{T}})}{\sqrt{N/r}}\mathrm{V}\end{gathered}\@add@centering$ (1)
@@
SparX introduces two flexible hyperparameters, namely S and M, making it easier and more efficient to extend the network deeper and wider. In addition, the incorporation of DMCA enables strong cross-layer feature interactions, giving rise to more powerful and robust representations.
@@
MSVMamba-T 4.6 33 82.8
SparX-Mamba-T 5.2 27 $\mathbf{83.5}$
@@
3.2 Network Architecture
Based on the proposed sparse cross-layer connections, we have developed a new Mamba-based vision backbone named SparX-Mamba, a 4-stage hierarchical architecture with three different variants: tiny, small, and base. In stage 1, we do not include any ganglion layers due to the high computational overhead when large-resolution inputs are used. That is, stage 1 contains only two normal layers. In stage 2, we designate the second layer as a ganglion layer, which has an intra-connection with its preceding layer. Since Stage 3 has more layers, we set different hyperparameters for the three variants. As listed in Table 1, due to the relatively small depth of the tiny model, we set S to 2 while S = 3 in both small and base models to avoid high computational cost in deeper models. In stage 4, we make an exception for tiny and small models by setting all layers as ganglion layers, as this stage processes the lowest resolution features and incurs a relatively low computational cost in these models. Regarding more complex base model, only the last layer is set as the ganglion layer to prevent excessive computational costs. Meanwhile, to facilitate communications across different stages, the first ganglion layer of each stage connects to a downsampled version of the final feature from the preceding stage. Furthermore, a cross-layer sliding window is only applied within each stage, to prevent high memory consumption due to excessive feature storage.